OpenStudy (masterchief):

Serious question, if you take something an multiply it by 1, you get the original number you started with. When you multiply 4 for instance by 1.5, it equals 6. What if you multiply 4 by 1 which equals 4, then by .5 which equals 2?

OpenStudy (anonymous):

Multiplication by 1 and then .5 is not the same as multiplication of 1.5, multiplication is commutative, and if that were true, then 1x.5 = 1.5, but it does not.

OpenStudy (mertsj):

4(1.5)=4(1+.5)=4(1)+4(.5)=4=2=6 Distributive Property

OpenStudy (anonymous):

However, multiplying 4 by 3 then .5 will result in 6 because 3 times .5 = 1.5